A has 20% more marbles than B, B has 20% more marbles than C. If A has 22 marbles more than C, how many marbles does B have?
Answer by josgarithmetic(39617)

Try to symbolize each statement.

a=b+0.2b
b=c+0.2c
a=c+22

Equating for a,
b+0.2b=c+22
c=b+0.2b-22

Substitute that for c into the equation for b.
b=c(1+0.2), then b=(b+0.2b-22)(1+0.2)
.
.
b=1.44b-26.4
b=60 marbles.